OpenStudy (anonymous):

If f(x) = e^xg(x), where g(0) = 5 and g'(0) = 3, find f '(0).

OpenStudy (anonymous):

by product rule (uv)' = u v' + v u' f'(x) = e^x g'(x) + g(x) e ^x f'(0) = e^0 3 + 5 e^0 = 3+5 = 8

OpenStudy (anonymous):

thank you!